OpenStudy (anonymous):

What is 3 root of 384? A. 3 3 root of 6 B. 4 3 root of 3 C. 4 3 root of 6 D. 6 3 root of 3

OpenStudy (anonymous):

A. \[3 \sqrt[3]{6}\]

OpenStudy (anonymous):

B. \[4 \sqrt[3]{3}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

C. \[4 \sqrt[3]{6}\]

OpenStudy (anonymous):

D. \[6 \sqrt[3]{3}\]

OpenStudy (anonymous):

there i translated it for all ya

OpenStudy (anonymous):

Do you have any idea?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

why is it C

OpenStudy (anonymous):

4*4*4*6=384

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

np tag me if u have any more questions

OpenStudy (anonymous):

ok